During 2017, Sandhill Furniture Company purchased wicker chairs. The manufacturer sells the chairs to Sandhill for a lump sum of $60,400. Three types of chairs are included in the carload. The three types and the estimated selling price for each are listed below.

Type

No. of Chairs

Estimated Selling
Price Each

Lounge chairs 400 $90
Armchairs 300 80
Straight chairs 700 50


During 2017, Sandhill sells 200 lounge chairs, 120 armchairs, and 130 straight chairs.

What is the amount of gross profit realized during 2017? What is the amount of inventory of unsold straight chairs on December 31, 2017? Round to 2 decimal places.

Gross profit realized during 2017?

Amount of inventory of unsold straight chairs?
Answer #1
Sandhill Furniture Company
Calculation of Cost per Chair
(A) (B) (C )=(A)*(B) (D )=(C )/95000 (E ) (F )=(D)*(E ) (G )=(F )/(A)
Number of chairs Selling Price per chair Total Selling Price Relative selling Price Total cost price Cost in percentage of relative selling price Cost per chair
Lounge Chairs 400 $                               90.00 $               36,000.00 $                               0.38 $          60,400.00 $                  22,888.42 $                       57.22
Arm Chairs 300 $                               80.00 $               24,000.00 $                               0.25 $          60,400.00 $                  15,258.95 $                       50.86
Straight Chairs 700 $                               50.00 $               35,000.00 $                               0.37 $          60,400.00 $                  22,252.63 $                       31.79
Total $               95,000.00
Calculation of Gross Profit realized
(A) (B) (C )=(A)*(B) (D ) (E )=(A)*(D) (F )=(C )-(E )
Number of chairs sold Selling Price per chair Total Selling Price Cost per chair Total cost price Gross Profit realized
Lounge Chairs 200 $                               90.00 $               18,000.00 $                             57.22 $          11,444.21 $                     6,555.79
Arm Chairs 120 $                               80.00 $                  9,600.00 $                             50.86 $            6,103.58 $                     3,496.42
Straight Chairs 130 $                               50.00 $                  6,500.00 $                             31.79 $            4,132.63 $                     2,367.37
Total $               34,100.00 $                  12,419.58
Gross Profit realized= $                      12,419.58
Amount of Inventory of unsold Straight chairs
(A) (B) (C )=(A)-(B) (D ) (C )*(D )
Number of units purchased Number of units sold Inventory(Units) Cost per chair Inventory cost
700 130 570 $                             31.79 18120